# Disentangled Representation Learning for Parametric Partial Differential Equations

### v1

@@ 3.1 Variational hierarchical neural operator as a multi-task solver
$L_{ELBO}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})-D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$
Putting everything together, we obtain the loss functional:
$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}\left[-E(\log(p(\bm{u}|\bm{f},\theta^{\eta})))-E_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})+D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})||p(\bm{z}^{\eta})\right)\right].$ (3.10)
The above formulation naturally lends itself to a hierarchical variational autoencoder (HVAE) architecture (Vahdat and Kautz, 2020). The encoder aims to obtain the posterior $q_{\mu_{z},\Sigma_{z}}(\bm{z}^{\eta}|\theta^{\eta})$, which estimates the inverse map $\mathcal{H}$:
$\hat{\bm{z}}\sim q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}^{\eta}|\theta^{\eta})\text{ .}$ (3.11)
Then, the first decoder, which processes the estimated latent variable $\hat{\bm{z}}$ and reconstructs the corresponding NO parameter $\hat{\theta}$:
$\hat{\theta}=\hat{g}(\hat{\bm{z}})\text{ .}$ (3.12)
@@
$\hat{\bm{u}}=\hat{\mathcal{G}}[\bm{f};\hat{\theta}]\text{ .}$ (3.13)
For this purpose, we employ the implicit Fourier neural operator (IFNO)***We also point out that the proposed strategy is generic and hence also applicable to other neural operators., which is a provably universal approximator for PDE solution operators with a relatively small number of trainable parameters (You et al., 2022b). To provide a universal $\hat{\mathcal{G}}$ for different domains, a multi-task learning strategy is adopted following Zhang et al. (2023), which applies task-wise adaptation only to the first layer of NOs. Specifically, an $L$-layer multi-task IFNO has the following form:
${\mathcal{G}}[\bm{f};{\theta}^{\eta}](\bm{x})={\mathcal{G}}[\bm{f};{\theta}_{P}^{\eta},\theta_{J},\theta_{Q}](\bm{x}):=\mathcal{Q}_{\theta_{Q}}\circ(\mathcal{J}_{\theta_{J}})^{L}\circ\mathcal{P}_{{\theta}^{\eta}_{P}}[\bm{f}](\bm{x})\text{ ,}$ (3.14)
where $\mathcal{P}$, $\mathcal{Q}$ are shallow-layer MLPs that map a low-dimensional vector into a high-dimensional vector and vice versa, parameterized by $\theta_{Q}$ and ${\theta}^{\eta}_{P}$, respectively. Each intermediate layer, $\mathcal{J}$, is constructed as a mimetic of a fixed-point iteration step and parameterized by $\theta_{J}$. Supported by the universal approximator analysis in Zhang et al. (2023), different PDEs share the common iterative ($\theta_{J}$) and projection ($\theta_{Q}$) parameters, with all the information about parameter $\bm{b}$ embedded in the task-wise lifting parameters $\theta^{\eta}_{P}$. Hence, the encoder and first decoder structure only need to encode and reconstruct $\theta^{\eta}_{P}$, since all other NO parameters are invariant to the change of $\bm{z}$. This construction substantially reduces the degree of freedoms in $\theta^{\eta}$, making the invertibility assumption in the next section feasible. With a slight abuse of notation, in the following contents we use $\theta^{\eta}$ to denote the task-wise parameter, i.e., the lifting layer parameter $\theta^{\eta}_{P}$, and estimate $\bm{z}^{\eta}$ from it.
@@ 3.2 Disentangling the underlying mechanism
The proposed DisentangO aims to identify and disentangle the components of the latent representation $\bm{z}$, which serves as an inverse PDE solver. However, to capture the true physical mechanism, a natural question arises: is it really possible to identify the latent variables of interest (i.e., $\bm{z}$) with only observational data $\mathcal{D}_{\rm tr}$? We now show that, by learning the model $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ that matches the true marginal data distribution in all domains, we can indeed achieve this identifiability for the generating process proposed in equation 3.5 and equation 3.6.
@@
We follow the data-generating process in equation 3.5 and equation 3.6, and make the following assumptions:
  • (Density Smoothness and Positivity): $p_{\theta}$ and $p_{\bm{z}}$ are both smooth and positive.
  • (Invertibility): The task-wise parameter $\theta$ can be generated by $\bm{z}$ through an invertible and smooth function $\mathcal{H}^{-1}$. Moreover, for each given $\bm{f}$, we denote $\mathcal{G}_{f}(\theta):=\mathcal{G}[\bm{f};\theta](\bm{x})$ as the operator mapping from $\mathbb{R}^{d_{\theta}}$ to $\mathcal{U}$. $\mathcal{G}_{f}$ is also one-to-one with respect to $\theta$.
Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve:
@@
We follow the data-generating process in equation 3.5 and equation 3.6, and assume that the Density Smoothness and Positivity condition as well as the Invertibility condition in Theorem 1 are both satisfied, together with the following additional assumptions:
  • (Conditional Independence): Conditioned on $\bm{b}$, each component of $\bm{z}$ is independent of each other: $\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i=1}^{d_{z}}\log p_{z_{i}|\bm{b}}(z_{i}|\bm{b})$.
  • (Linear independence) There exists $12d_{z}+$ 1 values of $\bm{b}$, such that the $2d_{z}$ vectors $w(\bm{z},\bm{b}^{j})-w(\bm{z},\bm{b}^{0})$ with $j=1,\cdots,2d_{z}$ are linearly independent. Here,
    $w(\bm{z},\bm{b}^{j}):=\left(\dfrac{\partial q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}},\cdots,\dfrac{\partial q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}},\dfrac{\partial^{2}q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}^{2}},\cdots,\dfrac{\partial^{2}q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}^{2}}\right),$ (3.16)
    with $q_{i}(z_{i},\bm{b}):=\log p_{z_{i}|\bm{b}}$.
Then, by learning $(p_{\hat{\bm{z}}},\hat{g},\hat{\mathcal{H}},\hat{\mathcal{G}})$ to achieve equation 3.15, $\bm{z}$ is component-wise identifiable.
@@
Intuitively, the linear independence condition in Theorem 2 indicates sufficient variability between physical systems (tasks). This is a common assumption in the nonlinear ICA literature for domain adaptation (Hyvarinen et al., 2019; Khemakhem et al., 2020; Kong et al., 2023). To our best knowledge, this is the first time the component-wise identifiability results are discussed in the context of multi-task neural operator learning.
3.3 A generic algorithm for different supervision scenarios
In this section, we discuss the practical utility of DisentangO under three scenarios:
  • (SC1: Unsupervised) No value or label is given for $\bm{b}^{\eta}$ for each task.
@@ 3.3 A generic algorithm for different supervision scenarios
  • (SC3: Supervised) The value of $\bm{b}^{\eta}$ is given.
The architecture is shown in Figure 1, with a pseudo algorithm provided in Algorithm 1 of Appendix.
To obtain the posterior $q$ in equation 3.11, we assume each latent variable satisfies a Gaussian distribution of distinct means and diagonal covariance, i.e., $q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}|\theta):=\mathcal{N}(\mu_{z}(\theta_{P}),\Sigma^{2}_{z}(\theta_{P}))$, then estimate its mean and covariance using an MLP. As a result, the KL-divergence term in the ELBO admits a closed form:
@@
In the supervised setting, we take $\bm{z}=\bm{b}$, and set $\mu_{z}(\theta_{P}^{\eta})=\bm{b}^{\eta}$, then:
@@
For the first decoder, the likelihood $p(\theta^{\eta}_{P}|\bm{z}^{\eta})$ is a factorized Gaussian with mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$, computed from another MLP. By taking as input Monte Carlo samples once for each $\bm{z}$, the reconstruction accuracy term can be approximated as:
$E_{q(\bm{z}|\theta_{P})}\log p(\theta_{P}|\bm{z})\approx-\sum_{i=1}^{d_{\theta}}\left(\dfrac{((\theta_{P})_{i}-(\mu_{\theta})_{i})^{2}}{2(\Sigma_{\theta})^{2}_{i}}+\log(\Sigma_{\theta})_{i}+c\right)\text{ ,}$
where ${d_{\theta}}$ is the dimension of $\theta_{P}$, and $c$ is a constant.
For the second decoder, we parameterize it as a multi-task NO in equation 3.14, and model the discrepancy between $\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})$ and the ground truth $\bm{u}(\bm{x}_{k})$ as an additive independent unbiased Gaussian random noise $\epsilon$, with
$\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k}))=\bm{u}_{j}^{\eta}(\bm{x}_{k})+\epsilon_{\eta,j,k}\text{ ,}\quad\epsilon_{\eta,j,k}\sim\mathcal{N}(0,\varpi^{2})\text{ ,}$
where $\varpi$ is a tunable hyperparameter. Then, with a uniform spatial discretization of size $\Delta x$ in a domain ${\Omega}\subset\mathbb{R}^{d_{u}}$, the log likelihood after eliminating the constant terms can be written as:
$\dfrac{1}{2\varpi^{2}\Delta x^{2d_{u}}}\sum_{\eta=1}^{S}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}](\bm{x}_{k})-\bm{u}_{j}^{\eta}(\bm{x}_{k})\right|\right|}_{L^{2}({\Omega})}^{2}.$
In our empirical tests in Section 4, for the unsupervised and semi-supervised scenarios we choose $q$ as a standard Gaussian distribution following the independence assumption of Theorem 2. Additionally, to avoid overparameterization, we take $\Sigma_{\theta}=\sigma^{2}_{\theta}\mathbf{I}$, with $\sigma_{\theta}$ being a tunable hyperparameter. In this case, the total objective can be further simplified as:
$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}\right)\text{ ,}$
@@
$L_{loss}=$ $\dfrac{1}{S}\sum_{\eta=1}^{S}\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}+\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}-\bm{b}^{\eta}\right|\right|}^{2}\right)\text{ ,}$
@@
$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}$ $\left(\beta_{d}\sum_{j=1}^{n_{train}}{\left|\left|\hat{\mathcal{G}}_{[\hat{\theta}^{\eta}_{P},\theta_{I},\theta_{Q}]}[\bm{g}^{\eta}_{j}]-\bm{u}^{\eta}_{j}\right|\right|}^{2}_{L^{2}({\Omega})}+{\left|\left|\hat{\theta}^{\eta}_{P}-\mu^{\eta}_{\theta}\right|\right|}^{2}\right.+\left.\beta_{KL}{\left|\left|\mu^{\eta}_{\bm{z}}\right|\right|}^{2}+\beta_{cls}L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))\right)$
where $\beta_{cls}$ is a tunable hyperparameter, and $L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))$ is the corresponding loss term depending on the provided information. For instance, when taking $c$ as the classification label of $\beta^{\eta}$, $L_{c}$ can be taken as the cross-entropy loss.
We note that the coefficient $\beta_{KL}$ is closely connected to the adjustable hyperparameter in $\beta-$VAE (Higgins et al., 2017). Taking a larger $\beta_{KL}$ encourages disentanglement (Locatello et al., 2019), while it also adds additional constraint on the implicit capacity of the latent bottleneck that results in information loss (Burgess et al., 2018). On the other hand, the data reconstruction term forces the latent factors to contribute to the “reconstruction” of the more complicated solution field in a global perspective, and hence increasing $\beta_{d}$ is anticipated to alleviate information loss. In our empirical studies, we demonstrate the interplay between these two tunable hyperparameters.

### v2

@@ 3.1 Variational hyper-neural operator as a multi-task solver
with
$\bm{f}\sim\mathbb{P}_{f},\,\bm{f}^{\eta}_{i,j}=\bm{f}^{\eta}_{i}(\bm{x}_{j}),\,\bm{u}^{\eta}_{i,j}=\bm{u}^{\eta}_{i}(\bm{x}_{j})+\epsilon_{\eta,i,j}=\mathcal{G}^{\dagger}[\bm{f}^{\eta}_{i};\bm{b}^{\eta}](\bm{x}_{j})+\epsilon_{\eta,i,j},\;\epsilon_{\eta,i,j}\sim\mathcal{N}(0,\varpi^{2}).$ (3.6)
The forward modeling objective is to learn a surrogate solution operator $\mathcal{G}[\cdot;\theta^{\eta}]$ for $\mathcal{G}^{\dagger}$, where $\theta^{\eta}$ denotes the task-specific NO parameters for the forward surrogate operator in the $\eta$-th task, associated with $\bm{b}^{\eta}$. Note that all tasks share the same NO architecture, $\mathcal{G}$, and the surrogate operator for each task depends on the physical system $\eta$ through the task-specific parameter $\theta^{\eta}$.
For inverse modeling, let $\mathcal{H}^{\dagger}$ denote the (unknown) inverse operator that maps function pairs $(\bm{u},\bm{f})$ to the underlying PDE parameter $\bm{b}$. In practice, identifying $\mathcal{G}^{\dagger}$is often impossible, as the available data of $(\bm{u},\bm{f})$ may not contain sufficient information to identify all features of $\bm{b}$. For instance, in a Dirichlet boundary condition problem, $\bm{u}^{\eta}_{i}(\bm{x})=0$ for all $\bm{x}\in\partial\Omega$, making $\bm{b}^{\eta}$ not learnable on $\partial\Omega$. A more realistic goal is therefore to recover the underlying mechanism of $\bm{b}$ in the space of identifiability, i.e., $\bm{z}$. Thus, our second objective is to construct an inverse map by estimating embedding $\tilde{\bm{z}}^{\eta}$ of the hidden parameter $\bm{b}^{\eta}$ from $\theta^{\eta}$:
@@
where $\Theta_{H}$ are trainable parameters of $\mathcal{H}$, and $\tilde{\bm{z}}^{\eta}\in\mathbb{R}^{d_{z}}$ denotes the learned latent variables that can be transformed to the ground-truth latent variable $\bm{z}^{\eta}$ via an invertible function $h$. The goal of disentanglement is to discover $\tilde{\bm{z}}$, together with the solution operator $\mathcal{G}$.
The first objective learns the forward operator $\mathcal{G}[\cdot;\theta^{\eta}]$ in a supervised fashion, in the form of function-to-function mappings, for all (hidden) parameters $\bm{b}^{\eta}$ in the range of interest; while the second objective learns the vector-to-vector mapping from $\theta^{\eta}$ to $\tilde{\bm{z}}^{\eta}$, which is unsupervised due to the hidden physics nature. We propose to employ a variational autoencoder (VAE) (Kingma & Welling, 2013) as the representation learning approach for the second objective, and a meta-learned NO architecture as a universal solution operator for the first objective***Note that tasks differ in the PDE coefficients (the physical parameters $b^{\eta}$) and correspondingly the NO parameter $\theta^{\eta}$, not in the NO architecture or its input-output structure.. The key is to pair these architectures as a hyper-neural operator, which simultaneously solves both forward and inverse problems. Although the proposed strategy is generic and thus applicable to other multi-task neural operator architectures, to provide a universal architecture of $\mathcal{G}$ for different tasks, we adopt a meta-learning strategy following the meta-learned neural operator (MetaNO) (Zhang et al., 2023). MetaNO is developed based on the implicit Fourier neural operator (IFNO), a PDE solution operator with a relatively small number of trainable parameters (You et al., 2022b). In MetaNO, task-wise adaptation is applied only to the trainable parameters in the first layer of the NO, whereas all other parameters are shared across tasks. For an $L$-layer MetaNO, we write:
[…]
where $\mathcal{P}$, $\mathcal{Q}$ are shallow-layer MLPs that map a low-dimensional vector to a high-dimensional vector and vice versa, parameterized by $\theta^{\eta}_{P}$ and $\theta_{Q}$, respectively. Each intermediate layer, $\mathcal{J}$, is constructed as a mimetic of a fixed-point iteration step and parameterized by $\theta_{J}$. Supported by the universal approximator analysis (Zhang et al., 2023), different PDEs share common iterative ($\theta_{J}$) and projection ($\theta_{Q}$) parameters, with all the information about parameter $\bm{b}$ encoded in the task-wise lifting parameters $\theta^{\eta}_{P}$. Hence, to provide an inverse map from the task-wise NO parameter $\theta^{\eta}$ to the key features of the PDE parameters, $\tilde{\bm{z}}^{\eta}$, one can construct $\mathcal{H}$ as the mapping from $\theta^{\eta}_{P}$ to $\tilde{\bm{z}}^{\eta}$ using a neural network architecture. In this work, we employ a standard MLP:
[…]
since all other NO parameters are invariant to the change of $\bm{b}$. This construction substantially reduces the degrees of freedom in $\theta^{\eta}$, making the invertibility assumption in the next section feasible. To simplify notation, we use $\theta^{\eta}$ to denote $\theta^{\eta}_{P}$ in the subsequent discussion. An overview of the forward and inverse operator architecture is provided in Figure 1.
We now define the learning objective. The overall objective is to maximize the log data likelihood:
$\max\mathbb{E}(\log p(\theta,\mathcal{D}))=\max\Big[\mathbb{E}\big(\log p(\bm{u}|\bm{f},\theta)+\log p(\theta|\bm{f})+\log p(\bm{f})\big)\Big].$
Note that $p(\bm{f})$ remains constant over different NO parameter $\theta$, and the assumption in eq. 3.6 yields $\log(p(\theta|\bm{f}))=\log(p(\theta))$. Additionally, the assumption in eq. 3.4 guarantees that $\theta$ is generated from the latent space over $\bm{z}$, hence $\log(p(\theta))=\log\int_{\bm{z}}p(\theta|\bm{z})p(\bm{z})d\bm{z}$. The overall objective then becomes:
$\max\mathbb{E}(\log(p(\theta,\mathcal{D})))=\max\bigg[\mathbb{E}(\log(p(\bm{u}|\bm{f},\theta)))+\mathbb{E}\Big(\log\int_{\bm{z}}p(\theta|\bm{z})p(\bm{z})d\bm{z}\Big)\bigg].$ (3.10)
@@
[…]
with $D_{KL}(\cdot||\cdot)$ denoting the KL divergence between two distributions. Putting everything together, we obtain the loss functional:
$L_{loss}=\frac{1}{S}\sum_{\eta=1}^{S}\bigg[-\mathbb{E}(\log(p(\bm{u}|\bm{f},\theta^{\eta})))-\mathbb{E}_{q(\bm{z}^{\eta}|\theta^{\eta})}\log p(\theta^{\eta}|\bm{z}^{\eta})+D_{KL}\left(q(\bm{z}^{\eta}|\theta^{\eta})\|p(\bm{z}^{\eta})\right)\bigg].$ (3.11)
The above formulation naturally lends itself to a hierarchical variational autoencoder (HVAE) architecture (Vahdat & Kautz, 2020). The encoder aims to obtain the posterior $q_{\mu_{z},\Sigma_{z}}(\bm{z}^{\eta}|\theta^{\eta})$ and provide the inverse map $\mathcal{H}$:
[…]
Note that although $\mathcal{H}$ represents a deterministic inverse mapping from the NO parameters to the latent variables, in equation 3.12 we approximate this mapping using a probabilistic encoder to account for uncertainty and enable variational inference. This follows standard practice in the VAE literature, where deterministic relationships are modeled probabilistically for tractable inference.
Then, the first decoder $\hat{g}$ processes the estimated latent variable $\hat{\bm{z}}$ and reconstructs the corresponding NO parameter $\hat{\theta}$:
[…]
@@
[…]
@@ 3.2 Disentangling the underlying mechanism
[…]
Assumption 1.
(Density Smoothness and Positivity) The probability density functions for $\theta$ and $\bm{z}$, which we denote as $p_{\theta}$ and $p_{\bm{z}}$, are both smooth and positive.
Assumption 2.
(Invertibility) The task-wise parameter $\theta$ can be generated by $\bm{z}$ through an invertible and smooth function $\mathcal{H}^{-1}$. Moreover, for each given $\bm{f}$, we denote $\mathcal{G}_{f}(\theta):=\mathcal{G}[\bm{f};\theta](\bm{x})$ as the operator mapping from $\mathbb{R}^{d_{\theta}}$ to $\mathcal{U}$. $\mathcal{G}_{f}$ is also one-to-one with respect to $\theta$.
Assumption 3.
(Conditional Independence) Conditioned on $\bm{b}$, each component of $\bm{z}$ is independent of each other: $\log p_{\bm{z}|\bm{b}}(\bm{z}|\bm{b})=\sum_{i=1}^{d_{z}}\log p_{z_{i}|\bm{b}}(z_{i}|\bm{b})$.
Assumption 4.
(Linear independence) There exists $2d_{z}+1$ values of $\bm{b}$, such that the $2d_{z}$ vectors $w(\bm{z},\bm{b}^{j})-w(\bm{z},\bm{b}^{0})$ with $j=1,\cdots,2d_{z}$ are linearly independent. Here,
$w(\bm{z},\bm{b}^{j}):=\Bigg(\dfrac{\partial q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}},\cdots,\dfrac{\partial q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}},\dfrac{\partial^{2}q_{1}(z_{1},\bm{b}^{j})}{\partial z_{1}^{2}},\cdots,\dfrac{\partial^{2}q_{d_{z}}(z_{d_{z}},\bm{b}^{j})}{\partial z_{d_{z}}^{2}}\Bigg),\vskip-5.78172pt$ (3.15)
with $q_{i}(z_{i},\bm{b}):=\log p_{z_{i}|\bm{b}}$.
@@
[…]
@@
[…]
@@
Discussion on assumptions. Intuitively, Assumptions 1-2 are required to guarantee that there exists a smooth and injective mapping between the ground-truth latent embedding $\bm{z}$ and the learned embedding $\hat{\bm{z}}$. Assumption 4 indicates sufficient variability across physical systems. This is a common assumption in the nonlinear ICA literature for domain adaptation (Hyvarinen et al., 2019; Khemakhem et al., 2020; Kong et al., 2023). Further discussions and validation on the assumptions are provided in Appendix A. To our best knowledge, it is the first time the component-wise identifiability is discussed in the context of multi-task neural operator learning.
3.3 A generic algorithm for various supervision cases
Although DisentangO is primarily designed for the challenging scenario of learning without supervision on $\bm{b}$ nor prior knowledge on the PDE model form in equation 3.1, its methodology is generic and readily can be extended to handle the scenarios with partial or full measurements of $\bm{b}$, which are common in classical inverse PDE benchmark problems. In this section, we discuss the practical utility of DisentangO under three scenarios:
  • (SC1: Supervised) The value of $\bm{b}^{\eta}$ is given.
@@ 3.3 A generic algorithm for various supervision cases
  • (SC3: Unsupervised) No value or label is given for $\bm{b}^{\eta}$ for each task.
A pseudo Algorithm 1 is provided in the Appendix.
To obtain the posterior $q$ in equation 3.12, we assume that each latent variable satisfies a Gaussian distribution of distinct means and diagonal covariance, i.e., $q_{\mu_{z},\Sigma_{z}}(\hat{\bm{z}}|\theta):=\mathcal{N}(\mu_{z}(\theta),\Sigma^{2}_{z}(\theta))$, then estimate its mean and covariance using an MLP. As a result, the KL-divergence term in the ELBO admits a closed form:
$\text{SC2/SC3:}\quad D_{KL}\left(q(\hat{\bm{z}}|\theta)\parallel p(\bm{z})\right)=\frac{1}{2}\sum_{i=1}^{d_{z}}\Big((\Sigma_{z})_{i}^{2}+(\mu_{z})^{2}_{i}-2\log((\Sigma_{z})_{i})-1\Big)\text{.}$ (3.17)
In the supervised setting, we take $\mu_{z}(\theta_{P}^{\eta})=\bm{b}^{\eta}$, and then the KL divergence term writes:
$\text{SC1:}\quad D_{KL}\left(q(\hat{\bm{z}}|\theta)\parallel p(\bm{z})\right)=\frac{1}{2}\sum_{i=1}^{d_{z}}\Big((\Sigma_{z})_{i}^{2}+(\mu_{z}-\bm{b})^{2}_{i}-2\log((\Sigma_{z})_{i})-1\Big)\text{.}\vskip-6.50403pt$ (3.18)
For the first decoder, the likelihood $p(\theta^{\eta}|\bm{z}^{\eta})$ is a factorized Gaussian with mean $\mu_{\theta}$ and covariance $\Sigma_{\theta}$, computed from another MLP. By taking as input Monte Carlo samples once for each $\bm{z}$, the reconstruction accuracy term can be approximated as:
@@
where $d_{\theta}$} is the dimension of $\theta$, and $c$ is a constant.
For the second decoder, we parameterize it as a multi-task NO in equation 3.8, and model the discrepancy between $\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}](\bm{x}_{k})$ and the ground truth $\bm{u}(\bm{x}_{k})$ as an additive independent unbiased Gaussian random noise $\epsilon$, with
$\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}](\bm{x}_{k})=\bm{u}_{j}^{\eta}(\bm{x}_{k})+\epsilon_{\eta,j,k}\text{, }\epsilon_{\eta,j,k}\sim\mathcal{N}(0,\varpi^{2})\text{,}$
where $\varpi$ is the standard deviation of the additive noise as defined in the data model. In practice, the observational noise is unknown, and we treat $\varpi$ as a tunable hyperparameter. Then, with a uniform spatial discretization of size $\Delta x$ in a domain ${\Omega}\subset\mathbb{R}^{d_{u}}$, the log likelihood after eliminating the constant terms can be written as:
@@
In our empirical tests in Section 4, for the unsupervised and semi-supervised scenarios, we choose $q$ as a standard Gaussian distribution following the independence assumption of Theorem 2. Additionally, to avoid overparameterization, we take $\Sigma_{\theta}=\sigma^{2}_{\theta}\mathbf{I}$, with $\sigma_{\theta}$ being a tunable hyperparameter. In this case, the total objective can be further simplified as:
[…]
@@
[…]
@@
$L_{loss}=\dfrac{1}{S}\sum_{\eta=1}^{S}\Bigg(\beta_{d}\sum_{j=1}^{n_{train}}\left\|\hat{\mathcal{G}}[\bm{f}^{\eta}_{j};\hat{\theta}^{\eta}]-\bm{u}^{\eta}_{j}\right\|^{2}_{L^{2}({\Omega})}+\left\|\hat{\theta}^{\eta}-\mu^{\eta}_{\theta}\right\|^{2}+\beta_{KL}\left\|\mu^{\eta}_{\bm{z}}\right\|^{2}+\beta_{cls}L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))\Bigg)$
with $\beta_{cls}$ a tunable parameter and $L_{c}(\bm{z}^{\eta},c(\beta^{\eta}))$ the corresponding loss term based on provided information (e.g., $L_{c}$ is the cross-entropy loss when taking $c$ as the classification label of $\beta^{\eta}$).
Note that $\beta_{KL}$ is closely connected to the adjustable hyperparameter in $\beta$-VAE (Higgins et al., 2017). Taking a larger $\beta_{KL}$ encourages disentanglement (Locatello et al., 2019), while it also adds additional constraint on the implicit capacity of the latent bottleneck resulting in information loss (Burgess et al., 2018). On the other hand, the data reconstruction term forces the latent factors to contribute to the “reconstruction” of the more complicated solution field in a global perspective, and hence increasing $\beta_{d}$ is anticipated to alleviate information loss. In our empirical study, we demonstrate the interplay between these two tunable hyperparameters.